# Model Predictive Control of the Neural Manifold

### v1

@@ Abstract
Neural manifolds are an attractive theoretical framework for characterizing the complex behaviors of neural populations. However, many of the tools for identifying these low-dimensional subspaces are correlational and provide limited insight into the underlying dynamics. The ability to precisely control this latent activity would allow researchers to investigate the structure and function of neural manifolds. Employing techniques from the field of optimal control, we simulate controlling the latent dynamics of a neural population using closed-loop, dynamically generated sensory inputs. Using a spiking neural network (SNN) as a model of a neural circuit, we find low-dimensional representations of both the network activity (the neural manifold) and a set of salient visual stimuli. With a data-driven latent dynamics model, we apply model predictive control (MPC) to provide anticipatory, optimal control over the trajectory of the circuit in a latent space. We are able to control the latent dynamics of the SNN to follow several reference trajectories despite observing only a subset of neurons and with a substantial amount of unknown noise injected into the network. These results provide a framework to experimentally test for causal relationships between manifold dynamics and other variables of interest such as organismal behavior and BCI performance.
keywords:
Neural Manifold, Model Predictive Control, Data-Driven Modeling, Optimal Control, Spiking Neural Network1 Introduction
Neural circuits are composed of large numbers of interconnected neurons whose activity depends on other neurons in the circuit. Because of these dependencies, simultaneously recorded populations of neurons usually exhibit high levels of correlation. Equivalently, most of the variance within the high-dimensional space corresponding to the firing rates of individual neurons is confined to a lower-dimensional subspace, or neural manifold [1]. Due to their relative simplicity, manifolds have become a popular framework for understanding the complex dynamics of large neural populations. In many different systems, activity on neural manifolds has been shown to correlate with salient features of stimuli, physical position, and internal cognitive states [2, 3, 4, 5]. However, the insight that manifolds can provide into the underlying dynamics and computational principles of the circuits remains a contentious question.
@@
where the latent dynamics on the manifold are defined by a function of the current state of the system , an external, time-varying stimulus that will be used for control, and an intrinsic, uncontrolled source of noise . The high dimensional measured neural activity is obtained with the observation function in equation (4). Interestingly, we show that the data-generating process (i.e., the dynamics of ) does not need to be the same as the latent dynamics model for the framework to be useful (see Methods) and that control of a highly nonlinear neural system is possible even when agnostic to the true structure of the neural manifold.
Given the latent dynamics specified in equation (3), the control problem is to find an external stimulus such that the time evolution follows a specified trajectory . The field of control theory provides a rich mathematical background to find system inputs to achieve desired system outputs. Many techniques exist with feedback (closed-loop) methods being particularly attractive due to their ability to correct for unknown perturbations to the system. Broadly speaking, feedback controllers can be categorized as being either reactive or anticipatory. Reactive controllers use present and past errors in state tracking (the difference between and ) to find a control signal . The classic PID controller is a popular implementation of a reactive controller due to its computational simplicity and strong performance. A notable issue with reactive controllers however is that one can only correct for errors in state tracking once they have been made. This can be unacceptable behavior if certain errors in state tracking are associated with pathological states. For example, suppose that a particular region of neural state space corresponded to epileptic firing. A reactive controller would only be able to respond after the system entered this region, at which point it may be much more difficult to re-establish control. In contrast, anticipatory controllers are designed to predict future errors, which can allow them to prevent the system from ever entering undesirable regions of state space. Model predictive control (MPC) is an anticipatory controller that uses a model of the system dynamics to make predictions on how present and future inputs will affect errors in state-tracking [14]. Additionally, MPC is a type of optimal controller because it attempts to find a control input sequence that minimizes a loss function of the form
@@
where is the value of the state at the current time step and is the loss associated with th time step (relative to ), which is a function of the state variable(s) and input(s) . The controller uses the dynamical model to simulate time steps into the future. Many types of loss functions are possible, but typically involve the state error and energy cost of the control signal. The constraints allow one to specify the dynamics of the system and to give lower and upper bounds for the state variables and inputs. More sophisticated versions of MPC allow for additional constraints where knowledge of any measurement or process noise can be incorporated [15]. Although MPC is only guaranteed to be globally optimal for linear systems with convex loss functions, it can also be used in many nonlinear systems [14, 16], in part because the controller can correct for model errors.
Only the first value of the sequence is used as input into the system, with the optimization performed again at the next time step. By repeatedly solving this optimization problem and only using the first value, the controller can correct for errors in system modeling and anticipate future changes in the desired state trajectory. This anticipation of dynamics can result in better controller performance compared to traditional reactive controllers such as PID [16].
There are two major issues when using MPC in practice that are relevant to control of neural dynamics. The first is that the optimization procedure is computationally expensive and can result in poor controller performance if the time steps between measurements are small [14]. Neural recordings are often of high dimension (e.g. extracellular recordings with high-density silicon probes) and evolve at fast time scales. The stimuli (corresponding to the input ) may also be of a high dimension, which will result in even more complexity in optimizing equation (5). We can now see the utility of modeling activity on the neural manifold as latent generative process. By doing so, we can solve the optimization problem in a low-dimensional state and reduce the computational complexity. If we used the descriptive approach as in equation (1), we would need to optimize in the original measurement dimension to force the activity on the manifold to follow a specified reference trajectory.
The second issue with applying MPC to neural systems that it requires a dynamical model of the system to be controlled [17]. Although there are many dynamical models rooted in biology for individual neurons, the putative latent dynamics of a neural manifold are an emergent property that is difficult to model from first principles. This requires us to take a data-driven approach, where unknown parts of the system can be modeled via function approximation and used to predict the time-evolution of the system in response to various inputs.
@@
These type of models are often referred to as forecasting models since the model predicts how the system will change across time. The goal is to find a set of parameters such that given some initial state value and some known input ,
@@
Data-driven approaches have been successfully applied to MPC problems in diverse fields [18, 19, 15, 20, 21], and the field is rapidly growing. In order for data-driven models to be useful for MPC applications in neuroscience, these models must be able to accurately predict the states to be controlled based only on observable state measurements, be agnostic to the number of hidden states, and generalize to a control scheme where control signals may be outside the training set. These challenges are not unique to neuroscience, but are still important to consider when selecting a data-driven approach to model the system dynamics. MPC has already been successfully applied in multiple areas of neuroscience research. At the individual cell level, simulated Hodgkin-Huxley neurons have been shown to be controllable via MPC using both biophysical [22, 23] and data-driven dynamics models [24, 25]. Additionally, both simulated and in vivo systems of neurons have been controlled with MPC using optogenetic stimulation as the control signal [26]. However, to the best of our knowledge, there has been no explicit attempt to apply these methods to activity on the neural manifold.
@@
Our goal for this study was to provide a framework for controlling the latent dynamics on a neural manifold with MPC. An artificial neural circuit was simulated using a spiking neural network (SNN) that was driven by images of handwritten digits . The activity of the network was measured in accordance with an extracellular recording experiment where the spike times of the SNN served as neural states . Modern extracellular probes are able to record from dozens to hundreds of neurons simultaneously in both anesthetized and awake animals, but this is only a small subset of the neurons in a typical local circuit. Thus, in this simulation we only used a subset of the simulated neurons to fit a latent dynamics model of the whole network. We also added a substantial amount of random synaptic noise to simulate the influence of uncontrolled spontaneous activity arising from unobserved neurons within and outside the local circuit.
@@
Using variational autoencoders (VAEs), we found low-dimensional representations of both the neural states and stimulus images. A latent dynamics model for the controller was fit using the lower-dimensional states and inputs (referred to as and respectively). We then found optimal latent inputs via MPC to force the latent states on the manifold to follow specified reference trajectories . We first controlled the latent dynamics of the SNN to stay at specified set points in the presence of intrinsic system noise. This result provided a proof of principle that MPC can control activity of the neural manifold in a simulation of a experiment using extracellular recording. We then show that the controller is able to force the latent states to follow multiple time-varying reference trajectories, with the optimized visual stimuli showing striking differences. The ability to control the system to follow distinct trajectories would allow for experimenters to see if there is a causal relationship between activity on the manifold and macro-scale behaviors. Finally, we examined the relationship between the proportion of measurable neurons and MPC performance.
@@ 2.1.1 Architecture
The activity of an artificial circuit evoked by an external visual stimulus was simulated with a SNN composed of three layers: sensory, reservoir, and output. Each neuron in the SNN was modeled with a recurrent leaky integrate-and-fire (rLIF) model, with the discrete time approximation
@@
Each neuron in the sensory layer received feedforward input in the form of a grayscale image reshaped into a 784 dimensional vector. This input served as the external stimulus that was the primary driver of SNN activity. Gaussian noise was also added to the feedforward inputs of each neuron in every layer to produce stochasticity in activity. This noise modeled the effects of natural variability in neural firing and the effects of unknown exogenous inputs. Thus the subthreshold activity of the three layers was given by the equations
where for each element in the feedforward input and is the stimulus image presented at the th time step. Note that incorporates both the membrane time constant and the time step of the discretization of the continuous LIF [27]. The time units for the simulation can be arbitrarily chosen by modifying this parameter. For simplicity, we use the units of milliseconds (ms) for each time step in all simulations which implicitly imposes a spiking refractory period of 1 ms. See Supplementary Information T1 for SNN training and rLIF hyperparameters.
@@ 2.1.2 Training the Network
In principle the feedforward and recurrent weights for each neuron could be randomly distributed, such as in reservoir computing [28]. However, to give the network dynamics that implement a useful computation [29], the weights were trained to perform a classification task. Using the Python package snntorch [27], the SNN was trained to accurately classify digits from the MNIST data set [30]. Training SNNs requires additional considerations compared to traditional artificial neural networks. The resetting of a neuron’s membrane voltage when it reaches the threshold produces a non-differentiable function [27] making training with gradient descent impossible. One solution to this problem is to use surrogate gradient descent [31], where the non-differentiable function is preserved in the forward pass of the network but is replaced with a sigmoid function during the backward pass. This results in a function differentiable everywhere and allows the network to be trained with valid gradients.
The MNIST data set is composed of handwritten images of the digits 0 through 9. Classification of the images was performed by associating the label of the image with a corresponding neuron in the output layer of the SNN (e.g., label 4 with neuron 4). For time step , the cross-entropy was given by
@@
where the cross-entropy loss at each time step is summed for some trial time length . This forces the neuron of the associated predicted class to have the highest firing rate compared to the other neurons in the output layer. For the given time window that the image is presented, the feedforward and recurrent weights in the SNN are updated using backpropagation through time (BPTT).
@@
The first half of the MNIST data set () was used for the training and validation of the SNN with an 80/20 split. Due to the stochastic nature of the noise added to every neuron in the AC, performance after training was assessed by presenting the stimuli 30 times to obtain an accuracy distribution. Accuracy on the training (n=28,000, %, %) and validation (, %, %) set were largely similar, indicating there was no overfitting to the training data. Although the accuracies were far below what would be considered competitive performance on a classification task, the purpose of training the SNN was to ensure the connections between the neurons were not random.
2.2 Dimensionality Reduction of Stimuli and Neural States
The dimensionalities of the MNIST digit stimuli and SNN neural activity were reduced using variational autoencoders (VAEs). The parameters for encoders and decoders were found by minimizing the loss function
where is a hyperparameter that scales the importance the of KL-divergence term in the loss function and is the batch size used for gradient descent. Because the latent variable is parameterized by a Gaussian, minimizing this loss function is equivalent to maximizing the evidence lower bound (ELBO) [32].
@@ 2.2 Dimensionality Reduction of Stimuli and Neural States
The VAE framework was chosen for two reasons. First, VAEs embed high-dimensional data into nonlinear low-dimensional subspaces. This allows for a flexible approach for finding the latent dynamics of the SNN and obtaining a more generalizable compression of the data compared to linear transformations [33]. Second, the use of KL-divergence promotes nearby values of in the latent space to be decoded into similar values in the original space of [34]. By doing so, this allows for easier interpolation between a set of training data points in latent space when constructing the dynamics model and finding the optimal latent inputs with MPC. If a basic autoencoder (i.e., ) was used instead of a VAE, nearby points in latent space would not be guaranteed to be similar in the original measurement space, which would obviously be detrimental to modeling dynamics in the latent space.

### v2

# Model Predictive Control on the Neural Manifold Authors: Christof Fehrman1 and C Daniel Meliza2,3

@@ Abstract
Neural manifolds are an attractive theoretical framework for characterizing the complex behaviors of neural populations. However, many of the tools for identifying these low-dimensional subspaces are correlational and provide limited insight into the underlying dynamics. The ability to precisely control the latent activity of a circuit would allow researchers to investigate the structure and function of neural manifolds. We simulate controlling the latent dynamics of a neural population using closed-loop, dynamically generated sensory inputs. Using a spiking neural network (SNN) as a model of a neural circuit, we find low-dimensional representations of both the network activity (the neural manifold) and a set of salient visual stimuli. The fields of classical and optimal control offer a range of methods to choose from for controlling dynamics on the neural manifold, which differ in performance, computational cost, and ease of implementation. Here, we focus on two commonly used control methods: proportional-integral-derivative (PID) control and model predictive control (MPC). PID is a computationally lightweight controller that is simple to implement. In contrast, MPC is a model-based, anticipatory controller with a much higher computational cost and engineering overhead. We evaluate both methods on trajectory-following tasks in latent space, under partial observability and in the presence of unknown noise. While both controllers in some cases were able to successfully control the latent dynamics on the neural manifold, MPC consistently produced more accurate control and required less hyperparameter tuning. These results demonstrate how MPC can be applied on the neural manifold using data-driven dynamics models, and provide a framework to experimentally test for causal relationships between manifold dynamics and external stimuli.
1 Introduction
Neural circuits are composed of large numbers of interconnected neurons whose activity depends on other neurons in the circuit. Because of these dependencies, simultaneously recorded populations of neurons usually exhibit high levels of correlation. Equivalently, most of the variance within the high-dimensional space corresponding to the firing rates of individual neurons is confined to a lower-dimensional subspace, or neural manifold [2]. Due to their relative simplicity, neural manifolds have become a popular framework for understanding the complex dynamics of large neural populations. In many different systems, activity on neural manifolds has been shown to correlate with salient features of stimuli, physical position, and internal cognitive states [3, 4, 5, 6]. However, the insight that manifolds can provide into the underlying dynamics and computational principles of the circuits remains a contentious question.
@@ 1 Introduction
where the latent dynamics on the manifold are defined by a function of the current state of the system , an external, time-varying stimulus that will be used for control, and an intrinsic, uncontrolled source of noise . The high dimensional measured neural activity is obtained with the observation function in equation (4). Interestingly, we show that the data-generating process (i.e., the dynamics of ) does not need to be the same as the latent dynamics model (LDM) for the framework to be useful (see Methods) and that control of a highly nonlinear neural system is possible even when agnostic to the true structure of the neural manifold.
1.1 Controlling System Dynamics
Given the latent dynamics specified in equation (3), the control problem is to find an external stimulus such that the time evolution follows a specified trajectory . The field of control theory provides a rich mathematical background to find system inputs to achieve desired system outputs. Many techniques exist with feedback (closed-loop) methods being particularly attractive due to their ability to correct for unknown perturbations to the system. Broadly speaking, feedback controllers can be categorized as being either reactive or anticipatory. Reactive controllers can use present and past errors in state tracking (the difference between and ) to find a control signal . Proportional-integral-derivative (PID) control is a popular implementation of a reactive controller due to its computational simplicity and strong performance. The control input is given by
where is the state error at time with gain matrices , , and . A discrete-time version of equation (5) is easily obtained by replacing the integral term with a sum and derivative term with finite differences. Although good performance can typically be guaranteed only for linear systems, PID is still a standard method for controlling nonlinear neural systems, such as in whole-cell voltage clamp experiments.
A notable issue with reactive controllers is that corrections can only be made for errors in state tracking once they have occurred. This can be unacceptable if certain errors in state tracking are associated with pathological states. For example, suppose that a particular region of neural state space corresponded to epileptic firing. A reactive controller would only be able to respond after the system entered this region, at which point it might be more difficult to re-establish control. In contrast, anticipatory controllers are designed to predict future errors, which can allow them to prevent the system from ever entering undesirable regions of state space. Model predictive control (MPC) is an anticipatory controller that uses a model of the system dynamics to make predictions on how present and future inputs will affect errors in state tracking [15]. Additionally, MPC is a type of optimal controller because it attempts to find a control input sequence that minimizes a loss function of the form
@@ 1.1 Controlling System Dynamics
where is the value of the state at the current time step and is the loss associated with th time step (relative to ), which is a function of either the state variable(s) and input(s) or both. The controller uses the dynamical model to simulate time steps into the future. Many types of loss functions are possible, but typically involve the state error and energy cost of the control signal. The constraints allow one to specify the dynamics of the system and to give lower and upper bounds for the state variables and inputs. More sophisticated versions of MPC allow for additional constraints where knowledge of any measurement or process noise can be incorporated [16]. Although MPC is only guaranteed to be globally optimal for linear systems with convex loss functions, it can also be used in many nonlinear systems [15, 17], in part because the controller can correct for model errors.
Only the first value of the sequence is used as input into the system, with the optimization performed again at the next time step. By repeatedly solving this optimization problem and only using the first value, the controller can correct for errors in system modeling and anticipate future changes in the desired state trajectory. This anticipation of dynamics can result in better controller performance compared to PID and other reactive controllers [17].
MPC has two major practical issues that are relevant to control of neural dynamics. The first is that the optimization procedure is computationally expensive and can result in poor controller performance if the time steps between measurements are small [15]. Neural recordings are often of high dimension (e.g., extracellular recordings with high-density silicon probes) and evolve at fast time scales. The stimuli (corresponding to the input ) may also be of a high dimension, which will result in even more complexity in optimizing equation (6). The utility of modeling activity on the neural manifold as a latent generative process is that we can reduce the computational complexity by solving the optimization problem in a lower-dimensional state space . If we used the descriptive approach as in equation (1), we would need to optimize in the original measurement dimension to force the activity on the manifold to follow a specified reference trajectory.
The second issue with applying MPC to neural systems that it requires a dynamical model of the system [18]. Although there are many dynamical models rooted in biology for individual neurons, the putative latent dynamics of a neural manifold are an emergent property that is difficult to model from first principles. This requires us to take a data-driven approach, where unknown parts of the system can be modeled via function approximation and used to predict the time-evolution of the system in response to various inputs.
@@
These type of models are often referred to as forecasting models, because the model predicts how the system will change across time. The goal is to find a set of parameters such that given some initial state value and some known input ,
@@
Data-driven approaches have been successfully applied to MPC problems in diverse fields [19, 20, 16, 21, 22], and the field is rapidly growing. In order for data-driven models to be useful for MPC applications in neuroscience, these models must be able to accurately predict the states to be controlled based only on observable state measurements, be agnostic to the number of hidden states, and generalize to a control scheme where control signals may be outside the training set. These challenges are not unique to neuroscience, but are still important to consider when selecting a data-driven approach to model the system dynamics. MPC has already been successfully applied in multiple areas of neuroscience. At the individual cell level, simulated Hodgkin-Huxley neurons have been shown to be controllable via MPC using both biophysical [23, 24] and data-driven dynamics models [25, 26]. Additionally, both simulated and in vivo systems of neurons have been controlled with LQR (a related optimal control method) using optogenetic stimulation as the control signal [27]. However, to the best of our knowledge, there has been no explicit attempt to apply these methods to activity on the neural manifold.
1.2 Study Overview
Our goals for this study are two-fold. First, we develop a framework for explicitly controlling the latent dynamics on a neural manifold. Second, we compare the performances of PID and MPC across several control tasks. Given that PID is computationally simple and does not require an explicit model of the system dynamics, the advantages of MPC must significantly outweigh the costs in order to advocate its implementation.
@@ 1.2 Study Overview
We simulated neural activity with an artificial circuit (AC) using a spiking neural network (SNN) that was driven by images of handwritten digits (). The activity of the network was measured as if in an extracellular recording experiment where the spike times of the SNN served as neural states (). Modern extracellular probes are able to record from dozens to thousands of neurons simultaneously in both anesthetized and awake animals, but this is only a small subset of the neurons in a typical local circuit. Thus, in this simulation we only used a subset of the simulated neurons to fit a LDM of the whole network. We also added random synaptic noise to simulate the influence of uncontrolled spontaneous activity arising from unobserved neurons within and outside the local circuit.
@@
Using variational autoencoders (VAEs), we found low-dimensional representations of both the neural states and stimulus images (referred to as and respectively) and fit a linear LDM using these latent representations. Then with either PID or MPC, we found latent inputs to force the latent states on the manifold to follow specified reference trajectories . Figure 1 illustrates the control loop diagram. We compared the two control strategies across two experimental conditions. First, we controlled the latent dynamics of the AC to stay at specified set points in the presence of intrinsic system noise. This result provides a proof of principle that both PID and MPC can control activity on the neural manifold. In the second experiment, we examined the generalizability of the PID and MPC controller hyperparameters from Experiment I. Our results indicated that even though MPC is more computationally complex than PID, MPC is superior across random AC subsets and proportions of measurable neurons. For a final example of neural manifold control, we used MPC to force the latent states to follow multiple time-varying reference trajectories, with the optimized visual stimuli showing striking differences. The ability to control the system to follow distinct trajectories would allow for experimenters to test for causal relationships between activity on the manifold and macro-scale behaviors.
@@ 2.1.1 Architecture
The activity of an AC evoked by an external visual stimulus was simulated with a SNN composed of three layers: sensory, reservoir, and output. Each neuron in the AC was modeled with a recurrent leaky integrate-and-fire (rLIF) model, with the discrete time approximation
@@
Each neuron in the sensory layer received feedforward input in the form of a grayscale image reshaped into a 784-dimensional vector. This input served as the external stimulus that was the primary driver of AC activity. Gaussian noise was also added to the feedforward inputs of each neuron in every layer to produce stochasticity in activity. This noise modeled the effects of natural variability in neural firing and the effects of unknown exogenous inputs. The subthreshold activity of the three layers was given by
where for each element in the feedforward input and is the stimulus image presented at the th time step. Note that incorporates both the membrane time constant and the time step of the discretization of the continuous LIF [28]. The time units for the simulation can be arbitrarily chosen by modifying this parameter. For simplicity, we use units of milliseconds (ms) for each time step in all simulations, which implicitly imposes a spiking refractory period of 1 ms. Additionally, we clamped membrane voltages to a minimum value of to prevent the membrane voltage from becoming unrealistically low. See Appendix 1 for AC training and rLIF hyperparameters.
@@ 2.1.2 Training the Network
In principle, the feedforward and recurrent weights for each neuron could be randomly distributed, such as in reservoir computing [29]. However, to give the network dynamics that implement a useful computation [30], the weights were trained to perform a classification task. Using the Python package snntorch [28], the AC was trained to classify digits from the MNIST data set [31]. Training SNNs requires additional considerations compared to traditional artificial neural networks. The resetting of a neuron’s membrane voltage when it reaches the threshold produces a non-differentiable function [28] making training with gradient descent impossible. One solution to this problem is to use surrogate gradient descent [32], where the non-differentiable function is preserved in the forward pass of the network but is replaced with a sigmoid function during the backward pass. This results in a function differentiable everywhere and allows the network to be trained using standard backpropagation methods.
The MNIST data set is composed of handwritten images of the digits 0 through 9. Classification of the images was performed by associating the label of the image with a corresponding neuron in the output layer of the AC (e.g., label 4 with neuron 4). For time step , the cross-entropy was given by
@@
where the cross-entropy loss at each time step is summed for some trial time length . This forces the neuron of the associated predicted class to have the highest firing rate compared to the other neurons in the output layer [28]. For the given time window that the image is presented, the feedforward and recurrent weights in the SNN are updated using backpropagation through time (BPTT).
@@
The AC was trained and assessed using half of the entire MNIST dataset (n = 35,000). This partition was split into training, validation, and testing sets using a 60/20/20 split. During training, early stopping was employed if the validation accuracy did not improve after 5 epochs. Due to the stochastic nature of the noise added to every neuron in the AC, performance after training was assessed by presenting the stimuli 30 times to obtain an accuracy distribution. Accuracy on the training (n = 21,000, M = 92.5%, SD = 0.001%) and validation (n = 7000, M = 90.1%, SD = 0.001%) set were largely similar, indicating there was little overfitting to the training data. Although the accuracies were below what would be considered competitive performance on a classification task, the purpose of training the AC was to ensure the connections between the neurons were not random.
2.2 Dimensionality Reduction of Stimulus Set
The high dimensionality of the stimulus inputs for the AC introduces a difficulty in the control problem. If the control input is found in the original measurement space, the values of would need to specified for each pixel value of the visual stimulus. This creates both a hyperparameter problem (the number of PID gains is equal to the number of pixels) and a computational one (MPC has to optimize over pixel space). A key insight is that the MNIST digits exhibit well-known low-dimensional structure, so we opted to find the control input in the latent stimulus space, which we denote . Once found, the AC can be stimulated by projecting back into the original stimulus dimension to obtain .
@@ 2.2 Dimensionality Reduction of Stimulus Set
A VAE was trained to find a low-dimensional representation of MNIST digit stimuli (sVAE). The VAE framework was chosen for two reasons. First, VAEs embed high-dimensional data into nonlinear low-dimensional subspaces. This allows for a flexible approach for finding a more generalizable compression of the data compared to linear transformations [33]. Second, the use of KL-divergence in the VAE cost function promotes nearby values of in the latent space to be decoded into similar values in the original stimulus space [34]. Intuitively, we want small changes in to produce small changes in . If a basic autoencoder was used instead of a VAE, nearby points in latent stimulus space would not be guaranteed to be similar in the original stimulus space, which would obviously be detrimental to modeling the latent dynamics and finding suitable control inputs.
The second half of the MNIST data set (n = 35,000) was used for training the sVAE and constructing a latent input sequence for latent dynamics identification. The data set was split into training, validation, and testing sets with a 70/15/15 split. During training, if the validation loss did not improve after 5 epochs, early stopping of sVAE training was implemented. The sVAE was trained via stochastic gradient descent in Pytorch using the Adam optimizer with the standard VAE loss function,
where is the size of the batch, is the sVAE decoder, and are the predicted parameters of the latent Gaussian, and is the latent stimulus dimension which was chosen to be . See Appendix 2 and 3 for architecture and training hyperparameters.
A sequence of latent inputs for LDM identification was constructed using the testing set of images. Discrete points in the latent sVAE space were obtained by running k-means clustering (n = 120) on the low-dimensional embedding of the validation images. These centers were evenly split (n = 40) into training, validation, and testing sets for LDM identification. See Figure 2A for latent embeddings of testing images and the corresponding centers. These discrete points were converted to a continuous time series by one of three methods: step function, fast interpolation, slow interpolation. The step function method took a sequence of centers and held each center constant for 500 time steps (ms). The fast and slow methods took the sequence and linearly interpolated values between each element in the sequence but at different time scales (200 ms for the fast and 1000 ms for the slow). The , and sequences were 36.80 seconds long (36,800 time steps). These three methods of interpolation were used to perturb the dynamics of the latent system at different timescales and frequencies. Figure 2B,C shows the latent input sequences and their sVAE decoded values.
2.3 Neural Manifold and Latent Dynamics Model Identification
The SNN was stimulated using the sVAE-decoded , and input sequences with the resulting spiking activity used to find the neural manifold and identify a LDM. For our initial experiment, the activity of a random sample of 20% of the neurons (n = 122) in the AC was recorded. The purpose of this random sampling was to mimic the incomplete information that would be obtained in a real recording, in which only a subset of neurons are observable.
where was chosen to be and . The neural manifold of the AC is the low-dimensional space parameterized by a latent encoding of this measured neural activity.